trung:

Given the investment flow I(t) = 2400 square root(t) (a) Calculate the total capital formation during the first 4 years. (b) Find an expression for the annual capital formation during the N th year and hence find the first year in which the annual capital formation exceeds $4000.

Narad:

(a)\[\int\limits_{0}^{4}2400\sqrt{t}dt\] (b)\[\int\limits_{n}^{n+1}2400\sqrt{t}dt\]